# Charged scalars at finite electric field and temperature in the optimized perturbation theory

### v1

@@ Abstract
We study the symmetry breaking and restoration behavior of a self-interacting charged scalar field theory under the influence of a constant electric field and finite temperature. Our study is performed in the context of the optimized perturbation theory. The dependence of the effective potential with constant electric fields is established by means of the bosonic propagators in the Schwinger proper-time method. Explicit analytical expressions for the electric and thermal contributions are found. Our results show a decreasing behavior for the vacuum expectation value as a function of the electric fields, which is strengthened by the temperature effect. A first-order phase transition that occurs at zero /weak electric fields changes to a second-order phase transition under strong electric fields. The critical temperature for the phase transition exhibited a very weak dependence on the electric field. Additionally, we computed the vacuum persistence probability rate for the interacting theory, finding a peak at the critical point. The maximum value of this rate at the critical point is found to be independent of the coupling constant but depended solely on the magnitude of the electric field.
@@ II The model and the OPT implementation
and where $D_{\mu}=\partial_{\mu}+iqA_{\mu}$, with $A_{\mu}$ the Abelian gauge field, here considered as an external background field. As usual, writing the complex field $\phi$ as $\phi=(\phi_{1}+i\phi_{2})/\sqrt{2}$. As usual, we can choose to shift the field around the $\phi_{1}$ direction, $\phi_{1}\to\phi_{1}+\varphi$, where $\varphi$ is a background field and which the effective potential is a function of. When $\varphi$ minimizes the potential, $\varphi\equiv\bar{\varphi}$, it becomes the vacuum expectation value (VEV) of the scalar field, $|\langle\phi_{1}\rangle|=\bar{\varphi}$ and $|\langle\phi_{2}\rangle|=0$. For the tree-level potential, we have from Eq. (2) that $\bar{\varphi}\equiv\varphi_{0}=\pm\sqrt{6m^{2}/\lambda}$. Hence, in the symmetry broken phase $\phi_{1}$ is associated with the Higgs field, while $\phi_{2}$ is the Goldstone field. Their Feynman propagators, written in terms of the VEV, is given by
@@
In terms of the interpolated model, the propagators given by Eqs. (3) and (4) now become
@@
The effective potential at first order in the OPT is Duarte:2011ph
@@
$\int_{p}D_{\Omega}(p)=-\frac{\Omega^{2}}{(4\pi)^{2}}\frac{1}{\epsilon}+X(T,\eta),$ (12)
@@ III The effective potential at finite temperature and in a constant electric field
where $F_{\mu\nu}$ is the electromagnetic field tensor. Using that the electric and magnetic fields are expressed by $E^{i}=F^{i0}$ and $B^{i}=-\sum_{j,k}\varepsilon^{ijk}F^{jk}/2$, we obtain
@@ IV Numerical Results
In this Section, we explore the numerical results of the model. Let us start by studying the effective potential Eq. (27) as a function of the background field and for different values of temperature and electric field. Our results show that there is a threshold value for the electric field below which the phase transition is first order, while above it the transition becomes second order. In Fig. 1 we have used a representative choice of model parameters, e.g., $\lambda=1$ and $m=0.1M$ (we express all dimensional quantities in terms of the regularization scale $M$). This is a choice of parameters where the results start becoming more sensitive to the effects of the electric field and one that allows us to compare the results in the absence of $E$ more easily.
@@
Despite the different transition patterns and changes in the electric field, the critical temperature for phase transition remains relatively constant, indicating a very weak dependence on the external electric field. This intriguing result is further explored below. Previous studies of the same model in the absence of electric fields ($E=0$Piccinelli:2013jua also found a first-order phase transition. In Ref. Piccinelli:2013jua , the authors investigated the phase transition behavior of the complex scalar field using the ring diagram approximation. Our results align with this finding, as we also observe a first-order phase transition up to the threshold value of the electric field within the context of the OPT. While not explicitly shown here, we have explored the phase transition for different parameter values and found that the threshold value for the electric field depends on these choices. For example, with $\lambda=2$ and $m=M$, we obtain $qE_{\rm threshold}\simeq 0.08M^{2}$, while for $\lambda=1$ and $m=M$ we find that $qE_{\rm threshold}\gtrsim 0.048M^{2}$. In general, increasing the self-coupling tends to increase the threshold value for the electric field at which the phase transition changes from first-order to second-order.
Figure 2: The VEV as a function of the temperature (panel a) and as a function of the electric field. The model parameters chosen are $\lambda=1$ and $m=0.1M$.
It is also relevant to look at the behavior of the VEV as a function of temperature and electric field. In Fig. 2 we show the normalized VEV $\bar{\varphi}/\varphi_{0}$, where $\bar{\varphi}\equiv\varphi(qE,T)$ and $\varphi_{0}^{2}=6m^{2}/\lambda$, which is shown as a function of the temperature and for different values of the external electric field (panel a) and as a function of the electric field and different values of the temperature (panel b). In Fig. 2(a), we can see that at $qE=0$ we have the usual symmetry restoration pattern, e.g. as previously observed in Duarte:2011ph , in which the $\bar{\varphi}/\varphi_{0}$ decreases as a function of $T$. Note also the discontinuity of the VEV close to the critical temperature, indicating a first-order phase transition, which persists up to the threshold value for the electric field discussed above. Above the threshold value for $E$, the VEV always decreases in a smooth way, vanishing at the critical point, as expected for a second-order phase transition Note also that for higher values of $qE$ the magnitude of $\varphi/\varphi_{0}$ at intermediate values of temperature in between $T=0$ and $T=T_{c}$ cause a more prominent change of the VEV. This behavior is also seen more clearly in Fig. 2(b), where we show the behavior of the VEV as a function of the electric field. In all cases, we find that the critical temperature, $T_{c}$, can be well approximately by
@@
Figure 3: The optimum OPT mass parameter $\bar{\eta}(qE,T)$ as a function of the temperature and for different values for the electric field.
@@ Appendix A Evaluation of pure electric part of $X(\eta,T=0,qE)$
$\ln\Gamma(x)=-\gamma_{E}x+\ln(x)+\sum_{k=1}^{\infty}\left[\frac{x}{k}-\ln\left(1+\frac{x}{k}\right)\right],$ (38)
@@
$\Im\left[\ln\Gamma\left(-iy+\frac{1}{2}\right)\right]=-\gamma_{E}y-\arctan(2y)$
$+$ $\sum_{k=1}^{\infty}\left[\frac{y}{k}-\arctan\left(\frac{-2y}{2k+1}\right)\right].$
@@
$\frac{-iqE}{8\pi^{2}}i\Im\left[\ln\Gamma\left(-iy+\frac{1}{2}\right)\right]=\frac{qE}{8\pi^{2}}\left\{-\gamma_{E}y-\arctan(2y)\right.$
$\left.+\sum_{k=1}^{\infty}\left[\frac{y}{k}-\arctan\left(\frac{-2y}{2k+1}\right)\right]\right\},$
@@ Appendix B Evaluation of pure electric part of $Y(\eta,T=0,qE)$
which can be easily verified from their general expressions, given by Eqs. (12) and (13). Therefore, it is straightforward to obtain $Y(\eta,qE,T=0)$ by
[…]
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\Re Y(qE,T=0,\Omega)=\frac{\Omega^{4}}{32\pi^{2}}\ln\left(\frac{2qE}{M^{2}}\right)$ (49)
@@ Appendix C Imaginary part of functions $X(\eta,T,qE)$ and $Y(\eta,T,qE)$
where $Y_{qE}(\eta,qE)$ is the pure electric part and $Y_{qE,T}(\eta,T,qE)$ is thermo-electric contribution, given by
$X_{qE,T}(\eta,T,qE)$ $=$ $\frac{(qE)^{2}}{8\pi^{2}}\int_{0}^{\infty}\frac{dt}{t\sin(t)}e^{-t\frac{\Omega^{2}}{qE}}$ (51)
@@
$=$ $\frac{(qE)^{2}}{8\pi^{2}}\int_{0}^{\infty}\frac{dt\cot(t)}{t\cos(t)}e^{-t\frac{\Omega^{2}}{qE}}$
@@
$\Im X_{qE,T}(\eta,T,qE)$ $=$ $\frac{(qE)^{2}}{8\pi^{2}}\int_{0}^{\infty}\frac{dt}{t\cos(t)}e^{-t\frac{\Omega^{2}}{qE}}$
@@
-- $\frac{\Omega^{2}}{16\pi^{2}}\ln\left(\frac{M^{2}}{-iqE}\right).$
@@
It is important to remark here that, in the context of electric fields, there is no clear consensus on the form or even the existence of one-loop thermal contributions for the imaginary part of the effective potential. These possible thermal effects will contribute to the vacuum persistence probability rate Loewe:1991mn ; Elmfors:1994fw ; Hallin:1994ad ; Ganguly:1995mi ; Ganguly:1998ys ; Gies:1998vt ; Gies:1999vb In particular, the worldline method has been used to derive these thermal contributions to the vacuum persistence probability rate Medina:2015qzc ; Gould:2017fve ; Korwar:2018euc . These studies have shown that the thermal contributions become relevant (compared to the $T=0$ result) at a threshold temperature $T_{CW}=\frac{eE}{2\Omega}$. Below $T_{CW}$ the thermal bath’s energy is insufficient to create a charged particle-antiparticle pair accelerated by the electric field over their Compton wavelength. In this regime, thermal effects are negligible, and charged particle-antiparticle pairs are produced only through quantum processes. All the numerical results presented in Sec. IV focus on the region where these possible thermal contributions to the effective potential and vacuum persistence probability rate can be ignored, i.e., we focus on the regime $T<T_{CW}$.
@@ Appendix D Independence of the critical temperature with the electric fields
$\frac{6m^{2}}{\lambda}-4X(T_{c},qE,\bar{\eta})\approx 0,$ (60)
@@
$X(T_{c},qE,\bar{\eta})$ $=$ $\frac{qE}{8\pi^{2}}\int_{0}^{\infty}\frac{dt}{t\sin(t)}\sum_{n=1}^{\infty}e^{-\frac{n^{2}qE}{4T_{c}^{2}|\tan(t)|}}.$
@@
$X(T_{c},qE,\eta)$ $=$ $\sum_{n=1}^{\infty}\frac{qE}{8\pi^{2}}\int_{0}^{\infty}\frac{d\tau}{\sqrt{1+\tau^{2}}}e^{-\frac{n^{2}qE}{4T_{c}^{2}}\tau}$
×$\times\frac{4}{\pi}\sum_{k=0}^{\infty}\left[\frac{(-1)^{k}(2k+1)}{(2k+1)^{2}-\left(\frac{\pi}{2}\tan^{-1}\tau\right)}\right].$

### v2

@@ Abstract
We study the symmetry breaking and restoration behavior of a self-interacting charged scalar field theory under the influence of a constant electric field and finite temperature. Our study is performed in the context of the optimized perturbation theory. The dependence of the effective potential with constant electric fields is established by means of the bosonic propagators in the Schwinger proper-time method. Explicit analytical expressions for the electric and thermal contributions are found. Our results show a very weak decreasing behavior of the vacuum expectation value as a function of the electric field, which is strengthened by the temperature effect. A first-order phase transition that occurs at zero /weak electric fields changes to a second-order phase transition under strong electric fields. The critical temperature for the phase transition exhibited a very weak dependence on the electric field. Additionally, we computed the vacuum persistence probability rate for the interacting theory, finding a peak at the critical point. The maximum value of this rate at the critical point is found to be independent of the coupling constant but depended solely on the magnitude of the electric field.
@@ II The model and the OPT implementation
and where $D_{\mu}=\partial_{\mu}+iqA_{\mu}$, with $A_{\mu}$ the Abelian gauge field, here considered as an external background field. As usual, writing the complex field $\phi$ as $\phi=(\phi_{1}+i\phi_{2})/\sqrt{2}$. As usual, we can choose to shift the field around the $\phi_{1}$ direction, $\phi_{1}\to\phi_{1}+\varphi$, where $\varphi$ is a background field and which the effective potential is a function of. Then, the symmetry breaking potential is now given by
$V(\varphi,\phi_{1},\phi_{2})$ $=$ $-\frac{m^{2}}{2}[(\phi_{1}+\varphi)^{2}+\phi_{2}^{2}]$ (3)
$+\frac{\lambda}{4!}[(\phi_{1}+\varphi)^{2}+\phi_{2}^{2}]^{2}.$
When $\varphi$ minimizes the potential, $\varphi\equiv\bar{\varphi}$, by applying
$\frac{\partial V(\phi_{1},\phi_{2},\varphi)}{\partial\phi_{1}}\Bigr{|}_{\phi_{1}=\bar{\varphi},\phi_{2}=0,\varphi=\bar{\varphi}}=0,$ (4)
it becomes the vacuum expectation value (VEV) of the scalar field, $|\langle\phi_{1}\rangle|=\bar{\varphi}$ and $|\langle\phi_{2}\rangle|=0$. For the tree-level potential, we have from Eq. (2) that $\bar{\varphi}\equiv\varphi_{0}=\pm\sqrt{6m^{2}/\lambda}$. Hence, in the symmetry broken phase $\phi_{1}$ is associated with the Higgs field, while $\phi_{2}$ is the Goldstone field. Their Feynman propagators, written in terms of the VEV, is given by
@@
We can derive the Feynman rules by the interpolated Lagrangian density in Eq. (7) by changing the vertex $-i\lambda$ by $-i\delta\lambda$; the new term $\delta\eta^{2}\phi_{i}^{2}/2$ is interpreted as a new vertex in the OPT formalism. In terms of the interpolated model, the propagators given by Eqs. (5) and (6) now become
@@
The vacuum diagrams shown in Fig. 1, represent the effective potential contributions to order $\delta$ in the OPT.
Figure 1: Vacuum Feynman diagrams that contribute to order $\delta$ to the effective potential. Solid lines corresponds to the propagator of $\phi_{1}$, and dashed lines corresponds to $\phi_{2}$. The black dots corresponds to $\delta\eta^{2}$ insertions, while the vacuum diagrams with crosses are represents the counter terms.
By inserting the propagators (10) and (11) in the vacuum diagrams shown in Fig. 1 and expanding again in $\delta$, the effective potential at first order in the OPT is obtained to be given by Duarte:2011ph
@@
$\int_{p}D_{\Omega}(p)=-\frac{\Omega^{2}}{(4\pi)^{2}}\frac{1}{\epsilon}+X(\eta),$ (14)
@@ III The effective potential at finite temperature and in a constant electric field
where $F_{\mu\nu}$ is the electromagnetic field tensor. The electric and magnetic fields are expressed by $E^{i}=F^{i0}$ and $B^{i}=-\sum_{j,k}\varepsilon^{ijk}F^{jk}/2$. Hence, we can set for definiteness the only nonzero fields to $i=3$ in both cases in order to have the bosonic propagator in a constant parallel magnetic and electric fields oriented in the $z$ direction, to obtain
@@ IV Numerical Results
In this Section, we explore the numerical results of the model. Let us start by studying the effective potential Eq. (29) as a function of the background field and for different values of temperature and electric field. Our results show that there is a threshold value for the electric field below which the phase transition is first order, while above it the transition becomes second order444 Note that from universality, on expects that the phase transition for a complex scalar field to be of second-order. The result obtained here can in principle be attributed to the order that the OPT approximation is carried out. This is however a point that needs further study. However, note that the transition is very weak first order, where the variation of $T_{c}$ between $E=0$ and when the transition becomes second-order, obtained by increasing $E$, is very small. Note also that a first-order phase transition for the complex scalar field was also obtained previously in Ref. Ayala:2004dx by using a different approximation method.. In Fig. 2 we have used a representative choice of model parameters, e.g., $\lambda=1$ and $m=0.1M$ (we express all dimensional quantities in terms of the regularization scale $M$). This is a choice of parameters where the results start becoming more sensitive to the effects of the electric field and one that allows us to compare the results in the absence of $E$ more easily.
@@
Despite the different transition patterns and changes in the electric field, the critical temperature for phase transition remains relatively constant, indicating a very weak dependence on the external electric field. This intriguing result is further explored below. Previous studies of the same model in the absence of electric fields ($E=0$Ayala:2004dx also found a first-order phase transition. In Ref. Ayala:2004dx , the authors investigated the phase transition behavior of the complex scalar field using the ring diagram approximation. Our results align with this finding, as we also observe a first-order phase transition up to the threshold value of the electric field within the context of the OPT. While not explicitly shown here, we have explored the phase transition for different parameter values and found that the threshold value for the electric field depends on these choices. For example, with $\lambda=2$ and $m=M$, we obtain $qE_{\rm threshold}\simeq 0.08M^{2}$, while for $\lambda=1$ and $m=M$ we find that $qE_{\rm threshold}\gtrsim 0.048M^{2}$. In general, increasing the self-coupling tends to increase the threshold value for the electric field at which the phase transition changes from first-order to second-order.
Figure 3: The VEV as a function of the temperature (panel a) and as a function of the electric field (panel b). The model parameters chosen are $\lambda=1$ and $m=0.1M$. The inset in panel (a) helps to see the behavior in the region close to $T=0$.
It is also relevant to look at the behavior of the VEV as a function of temperature and electric field. In Fig. 3 we show the normalized VEV $\bar{\varphi}/\varphi_{0}$, where $\bar{\varphi}\equiv\varphi(qE,T)$ and $\varphi_{0}^{2}=6m^{2}/\lambda$, which is shown as a function of the temperature and for different values of the external electric field (panel a) and as a function of the electric field and different values of the temperature (panel b). In Fig. 3(a), we can see that at $qE=0$ we have the usual symmetry restoration pattern, e.g. as previously observed in Duarte:2011ph , in which the $\bar{\varphi}/\varphi_{0}$ decreases as a function of $T$. Note also the discontinuity of the VEV close to the critical temperature, indicating a first-order phase transition, which persists up to the threshold value for the electric field discussed above. At finite values of electric field considered, we can see that the different curves of $\bar{\varphi}/\varphi_{0}$ almost coincide at $T=0$, indicating a very-weak electric field dependence of the VEV at low temperatures. Above the threshold value for $E$, the VEV always decreases in a smooth way, vanishing at the critical point, as expected for a second-order phase transition Note also that for higher values of $qE$ the magnitude of $\varphi/\varphi_{0}$ at intermediate values of temperature in between $T=0$ and $T=T_{c}$ cause a more prominent change of the VEV. This behavior is also seen more clearly in Fig. 3(b), where we show the behavior of the VEV as a function of the electric field. In all cases, we find that the critical temperature, $T_{c}$, can be well approximately by
@@
Next, in Fig. 4 we explore the critical exponent, $\beta$, by performing a linear fitting of $\ln\left(\frac{\varphi}{\varphi_{0}}\right)$ as a function of $\ln\left(\frac{T-T_{c}}{T_{c}}\right)^{\beta}$. Due to the scale of Fig. 3(a), it may be difficult to determine whether the second-order phase transition occurs with the critical exponent $\beta=1/2$, as expected from mean-field theory. However, the linear fit indeed shows $\beta\approx 0.5$, which is in good agreement with universality class arguments. We have numerically tested different values of the electric fields, and all results agree well with the previous statement.
Figure 4: Linear fitting of $\ln\left(\frac{\varphi}{\varphi_{0}}\right)$ as a function of $\ln\left(\frac{T-T_{c}}{T_{c}}\right)$ for different values of electric fields $qE$.
Figure 5: The optimum OPT mass parameter $\bar{\eta}(qE,T)$ as a function of the temperature and for different values for the electric field. The model parameters chosen are $\lambda=1$ and $m=0.1M$.
@@ Appendix A Evaluation of pure electric part of $X(\eta,T=0,qE)$
$\ln\Gamma(x)=-\gamma_{E}x-\ln(x)+\sum_{k=1}^{\infty}\left[\frac{x}{k}-\ln\left(1+\frac{x}{k}\right)\right],$ (40)
@@
$\Im\left[\ln\Gamma\left(iy+\frac{1}{2}\right)\right]=-\gamma_{E}y-\arctan(2y)$
$+$ $\sum_{k=1}^{\infty}\left[\frac{y}{k}-\arctan\left(\frac{2y}{2k+1}\right)\right].$
@@
$\frac{-iqE}{8\pi^{2}}i\Im\left[\ln\Gamma\left(iy+\frac{1}{2}\right)\right]=\frac{qE}{8\pi^{2}}\left\{-\gamma_{E}y-\arctan(2y)\right.$
$\left.+\sum_{k=1}^{\infty}\left[\frac{y}{k}-\arctan\left(\frac{2y}{2k+1}\right)\right]\right\},$
@@ Appendix B Evaluation of pure electric part of $Y(\eta,T=0,qE)$
which can be easily verified from their general expressions, given by Eqs. (14) and (15). Therefore, it is straightforward to obtain $Y(\eta,T=0,qE)$ by
$\int d\Omega^{2}X(\eta,T=0,qE)=Y(\eta,T=0,qE),$ (50)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\Re Y(\eta,T=0,qE)=\frac{\Omega^{4}}{32\pi^{2}}\ln\left(\frac{2qE}{M^{2}}\right)$ (51)
@@ Appendix C Imaginary part of functions $X(\eta,T,qE)$ and $Y(\eta,T,qE)$
where $X_{qE}(\eta,qE)$ is the pure electric part and $X_{qE,T}(\eta,T,qE)$ is thermo-electric contribution, given by
$X_{qE,T}(\eta,T,qE)$ $=$ $\frac{qE}{8\pi^{2}}\int_{0}^{\infty}\frac{dt}{t\sin(t)}e^{-t\frac{\Omega^{2}}{qE}}$ (53)
@@
$=$ $\frac{qE}{8\pi^{2}}\int_{0}^{\infty}\frac{dt\cot(t)}{t\cos(t)}e^{-t\frac{\Omega^{2}}{qE}}$
@@
$\Im X_{qE,T}(\eta,T,qE)$ $=$ $\frac{qE}{8\pi^{2}}\int_{0}^{\infty}\frac{dt}{t\cos(t)}e^{-t\frac{\Omega^{2}}{qE}}$
@@
-- $\frac{\Omega^{2}}{16\pi^{2}}\ln\left(\frac{M^{2}}{-2iqE}\right).$
@@
It is important to remark here that, in the context of electric fields, there is no clear consensus on the form or even the existence of one-loop thermal contributions for the imaginary part of the effective potential. These possible thermal effects will contribute to the vacuum persistence probability rate Loewe:1991mn ; Elmfors:1994fw ; Hallin:1994ad ; Ganguly:1995mi ; Ganguly:1998ys ; Gies:1998vt ; Gies:1999vb . In particular, the worldline method has been used to derive these thermal contributions to the vacuum persistence probability rate Medina:2015qzc ; Gould:2017fve ; Korwar:2018euc . These studies have shown that the thermal contributions become relevant (compared to the $T=0$ result) at a threshold temperature $T_{CW}=\frac{eE}{2\Omega}$. Below $T_{CW}$ the thermal bath’s energy is insufficient to create a charged particle-antiparticle pair accelerated by the electric field over their Compton wavelength. In this regime, thermal effects are negligible, and charged particle-antiparticle pairs are produced only through quantum processes. All the numerical results presented in Sec. IV focus on the region where these possible thermal contributions to the effective potential and vacuum persistence probability rate can be ignored, i.e., we focus on the regime $T<T_{CW}$.
@@ Appendix D Independence of the critical temperature with the electric fields
$\frac{6m^{2}}{\lambda}-4X(\bar{\eta},T_{c},qE)\approx 0,$ (62)
@@
$X(\bar{\eta},T_{c},qE)$ $=$ $\frac{qE}{8\pi^{2}}\int_{0}^{\infty}\frac{dt}{t\sin(t)}\sum_{n=1}^{\infty}e^{-\frac{n^{2}qE}{4T_{c}^{2}|\tan(t)|}}.$
@@
$X(\bar{\eta},T_{c},qE)$ $=$ $\sum_{n=1}^{\infty}\frac{qE}{8\pi^{2}}\int_{0}^{\infty}\frac{d\tau}{\sqrt{1+\tau^{2}}}e^{-\frac{n^{2}qE}{4T_{c}^{2}}\tau}$
×$\times\frac{4}{\pi}\sum_{k=0}^{\infty}\left[\frac{(-1)^{k}(2k+1)}{(2k+1)^{2}-\left(\frac{2}{\pi}\tan^{-1}\tau\right)^{2}}\right].$